OpenStudy (anonymous):

ASAP HELP NEED A GOOD GRADE PLEASE ANYONE FROM CONNECTIONS AND DOING THE UNIT 5 TEST LESSON 11 PLEASE HELP!!!! 1. what is a polynomial function in standard form with zeroes 1,2,-2, and -3 A. x^4+2x^3+7x^2-8x+12 B. x^4+2x^3-7x^2-8x+12 C. x^4+2x^3-7x^2+8x+12 D. x^4+2x^3+7x^2+8x+12 2. what are the zeroes of the function? what are their multiplicity f(x)=5x^3-5x^2-30x A. the numbers 3,-2, and 0 are zeroes of multiplicity 1 B. the numbers 3,-2, and 0 are zeroes of multiplicity 2 C. the numbers -3,2, and 0 are zeroes of multiplicity 1 D. the numbers -3,2, and 0 are zeroes of multiplicity 2

OpenStudy (anonymous):

3. what are the real or complex solutions of the polynomial equation? x^2-216=0 A. 3,3+3i square root (3),3-3i square root (3) B 3,-3 + 3i square root (3), -3-3i square root (3) C. 6,3 + 3i square root (3), -3+3i square root (3) D. 6,-3 3i square root (3), -3-3i square root (3) 4. which is a factored form of 27x^3+64 A. (3x+4)(9x^2-24+16) B. (3x+4)(9x^2+24+16 C. (3x+4)(9x^2-12x+16) D. (3x+4)(9x^2+12x-16) 5. divid x^3-2x^210x+136=0 A. -3+-5i,-4 B. 3+-5i,-4 C. -3+- i,4 D. 3 +- i , 4 6. one zero of f*x)=x^3+2x^2-5x-6 is 2. what are the other zeroes of the function? A. 1 and 3 B. -1 and -3 C. 2 and -3 D. 2 and 3 7. what is the second term of (s+v)^5 A. 25s^4 B. 25s^4v C. 5s^4v D. 125s^4v 8. which of the following quartic functions has x=-2 and x=-3 as its only two real zeroes? A. x^4-5x^3+7x^2-5x+6 B. x^4+5x^3+7x^2+5x+6 C. x^4+5x^3+7x^2-5x-6 D. x^4-5x^3+7x^2+5x-6